### v2

# Language Arithmetics: Towards Systematic Language Neuron Identification and Manipulation

Abstract
Large language models (LLMs) exhibit strong multilingual abilities, yet the neural mechanisms behind language-specific processing remain unclear. We analyze language-specific neurons in Llama-3.1-8B, Mistral-Nemo-12B, and Aya-Expanse-8B & 32B across 21 typologically diverse languages, identifying neurons that control language behavior. Using the Language Activation Probability Entropy (LAPE) method, we show that these neurons cluster in deeper layers, with non-Latin scripts showing greater specialization. Related languages share overlapping neurons, reflecting internal representations of linguistic proximity.
Through language arithmetics, i.e. systematic activation addition and multiplication, we steer models to deactivate unwanted languages and activate desired ones, outperforming simpler replacement approaches. These interventions effectively guide behavior across five multilingual tasks: language forcing, translation, QA, comprehension, and NLI. Manipulation is more successful for high-resource languages, while typological similarity improves effectiveness. We also demonstrate that cross-lingual neuron steering enhances downstream performance and reveal internal "fallback" mechanisms for language selection when neurons are progressively deactivated. Our code is made publicly available at https://github.com/d-gurgurov/Language-Neurons-Manipulation.
Language Arithmetics: Towards Systematic Language Neuron Identification and Manipulation
Daniil Gurgurov1,2  Katharina Trinley1  Yusser Al Ghussin1,2 Tanja Bäumel1,2,3Josef van Genabith1,2Simon Ostermann1,2,3 1Saarland University 2German Research Center for AI (DFKI) 3Centre for European Research in Trusted AI (CERTAIN) daniil.gurgurov@dfki.de
1 Introduction
The emergence of large language models (LLMs) with impressive multilingual capabilities has raised fundamental questions about how these systems internally represent and process different languages Wendler et al. (2024); Zhao et al. (2024). While models like Llama-3 Grattafiori et al. (2024) and Gemma-3 Team et al. (2024) perform well across dozens of languages despite limited multilingual training data, the neural mechanisms underlying this competence, such as the emergence of language-specific spaces, cross-lingual feature sharing, and internal language routing, are not fully understood. Understanding these mechanisms is crucial not only for advancing theoretical insights into multilingual representation learning, but also for building more controllable and interpretable language technologies Amodei et al. (2016); Gabriel (2020); Singh et al. (2024).
Figure 1: Success rates of language forcing when deactivating neurons for the input language and activating those of a target language for Llama-3.1-8B.
Recent studies have begun to explore how multilingual LLMs process language internally, showing that specific neurons may specialize in particular languages (Tang et al., 2024; Kojima et al., 2024; Zhao et al., 2024; Mondal et al., 2025). However, these works often focus on narrow language sets or specific architectures, leaving open questions about how language-specific processing scales across typologically diverse languages and how these insights can support model control and enhancement.
In this work, we present the first large-scale investigation of language-specific neuron identification and manipulation in the Llama-3.1-8B Grattafiori et al. (2024), Mistral-Nemo-12B AI (2024), and Aya-Expanse-8B & 32B Dang et al. (2024) models across 21 typologically diverse languages. We examine the distribution of language-sensitive neurons, their cross-linguistic overlap, and the evolution of output language distributions across layers. Using the Language Activation Probability Entropy (LAPE) method (Tang et al., 2024), we identify neurons with strong language preferences and report several key findings. Unlike previously tested models such as Llama-2 Touvron et al. (2023), BLOOM Le Scao et al. (2023), and Phi-2 Javaheripi et al. (2023), the models we study exhibit concentrated language-specific activity in their deeper layers. Analyzing output language distributions with logit lens Wendler et al. (2024) across layers reveals that language generation predominantly occurs in the final layers, aligning with the concentration of language-sensitive neurons. Non-Latin script languages show both a higher number of specialized neurons and less cross-language overlap, while typologically related Germanic and Romance languages share more neurons–reflecting linguistic proximity and possibly orthographic similarities in their internal representations.
Building on these findings, we establish simple additive and multiplicative neuron intervention techniques, which we call language arithmetics [^1] The naming follows conceptually similar works on task and prompt arithmetics Ilharco et al. (2023); Belanec et al. (2024)., to test the causal role of language-specific neurons in controlling language use while preserving internal representations. Unlike prior approaches that rely on activation replacement Tang et al. (2024); Kojima et al. (2024), our approach adds target-language patterns to hidden states, while simultaneously multiplying unwanted language patterns with 0. We demonstrate that language arithmetics outperforms both replacement-based methods Tang et al. (2024); Kojima et al. (2024) and the widely used DiffMean Marks and Tegmark (2023) approach on a novel language forcing task, where models are expected to answer questions in a specific target language without being explicitly prompted, with especially strong results for languages better represented in pretraining. We further evaluate additive language arithmetic on four downstream tasks–machine translation, question answering, natural language inference, and machine comprehension–and show that targeted neuron activation improves performance without any task-specific fine-tuning, contrasting with the findings of Mondal et al. (2025). For example, our intervention improves translation scores by up to 10% and enables precise cross-lingual transfer. Additionally, we find that when dominant language signals are suppressed, models fall back to the next most probable language, suggesting the presence of internal "fallback" mechanisms for language selection.
Overall, our work (1) reveals new patterns of language-specific neuron specialization and layer-wise processing, (2) applies a minimally invasive steering technique to test their causal influence, and (3) demonstrates practical benefits for multilingual model control and performance.
2 Related Work
Recent research has focused on uncovering language-specific mechanisms in LLMs. Several works have proposed methods to identify and intervene on neurons specialized for particular languages.
Tang et al. (2024) introduce the Language Activation Probability Entropy (LAPE) method, demonstrating that certain neurons are critical to multilingual capacity and can be manipulated to control language behavior via activation (setting neurons to their average values) or deactivation (zeroing them out). Kojima et al. (2024), building on Cuadros et al. (2022), similarly identify neurons selectively active for one language and inactive for others, and show that setting these to their median activation can shift generation language. Tan et al. (2024) further propose a frequency-based method to rank FFN neurons by activation counts on language-specific inputs, and show that fine-tuning these neurons improves downstream tasks such as machine translation. Expanding beyond FFNs, Zhao et al. (2024) introduce Parallel Language-specific Neuron Detection (PLND), which identifies language neurons in both attention and FFN modules. They show that deactivating these neurons reduces performance in the corresponding language, and that fine-tuning them on limited language data enhances multilingual ability.
Meanwhile, Deng et al. (2025) highlight the problem of superposition–where neurons encode multiple concepts (Elhage et al., 2022)–and instead use sparse autoencoders (SAE) (Cunningham et al., 2023) to extract latent dimensions linked to language identity. Ablating these features degrades performance, and they use them to guide steering vector construction for better control. Following this, Chou et al. (2025) show that modifying a single SAE feature can steer generation language with high accuracy and semantic preservation, especially in mid-to-late layers. Similarly, Andrylie et al. (2025) introduce SAE-LAPE to identify interpretable language-specific features in the sparse weight space, enabling both language identification and control.
Our work extends this line of research in two dimensions: (1) We analyze the internal structure of LLMs without relying on auxiliary models such as SAEs, and (2) we look at a wider range of more diverse languages. We conduct a comprehensive study of language-specific neurons in LLaMA-3.1-8B Grattafiori et al. (2024), Mistral-Nemo-12B AI (2024), and Aya-Expanse-8B & 32B Dang et al. (2024) across 21 languages using the LAPE method (Tang et al., 2024), systematically investigating both neuron identification and manipulation to assess their effectiveness for multilingual control and downstream performance improvement.
3 Language Neuron Identification
3.1 Identification Method
Language Activation Probability Entropy (LAPE) Tang et al. (2024) identifies language-specific neurons within LLMs by analyzing activation patterns across different languages in the FFN modules of a transformer-based language model. For each neuron, it measures how often that neuron activates when processing inputs in various languages. These activation frequencies are normalized into a probability distribution over languages, and entropy Shannon (2001) is used to quantify how language-specific each neuron’s behavior is.
Intuitively, neurons with low LAPE values are considered language-specific since their activation probabilities are concentrated on one or two languages, showing minimal activity for others. We select the bottom K% (K={1..5}) of neurons by LAPE score after filtering out neurons with weak overall activity: neurons are excluded if none of their language activations exceed the 95th percentile of all activations (filter rate). To ensure neurons exhibit clear language preferences, they further get refined by verifying that their activation probability for a given language surpasses the 95th percentile (activation threshold). These thresholds help eliminate noisy or non-informative neurons and retain only those with high, selective activation characteristics.
3.2 Per-layer Output Language
We additionally analyze how language identity emerges across layers by computing three key statistics, using a method referred to as logit lens Nostalgebraist (2020), similar to Wendler et al. (2024). Specifically, we apply a FastText classifier Joulin et al. (2016) to determine the language of the model’s output at each layer. First, we track the probability of generating the correct target language, revealing the depth at which language-specific behavior becomes prominent on the output level. Second, we measure the probability assigned to English regardless of the input, assessing the extent of English interference across layers. Third, we compute the entropy of the output language distribution to quantify the diversity and confidence of the model’s predictions.
3.3 Models and Data
We use the Llama-3.1-8B base model Grattafiori et al. (2024) for all experiments. Notably, Llama-3.1 is a decoder-only model that was neither explicitly trained for multilingual tasks nor instruction-tuned. Its pre-training data includes approximately 5% non-English content spanning over 30 languages. We also include Mistral-Nemo Mistral AI Team (2024), a slightly larger 12-B-parameter base model, featuring strong multilingual, reasoning, and coding performance in its size class. Finally, we evaluate the Aya Expanse family Dang et al. (2024): Aya-Expanse-8B, an open-weight, instruction-tuned multilingual model optimized via data arbitrage, preference training, and model-merging; and its larger sibling, Aya-Expanse-32B–both supporting 23 languages and offering state-of-the-art multilingual performance.
For neuron identification, we use the CulturaX corpus Nguyen et al. (2023), a large-scale multilingual dataset comprising over 6.3 trillion tokens across 167 languages. The data is sourced from Common Crawl Wenzek et al. (2019) and Wikipedia Foundation and has undergone extensive cleaning and language identification to ensure high quality. To compute language-specific activations and obtain LAPE values, we truncate the dataset for each language to 500MB due to efficiency reasons. For per-layer output experiments, we use 6 questions that are available in all languages, as described in Section 4.2.
Figure 2: Neuron overlap between languages and language families in Llama-3.1, based on the top 1% of neurons identified as language-specific. Diagonals show counts per language; off-diagonals show overlaps. Asterisks mark non-Latin script languages.
3.4 Results
Language Neuron Distribution: Figure 3 and Appendix E show total neuron activation distributions with minimal early-layer activity and pronounced peaks in later layers across all 21 languages. Interestingly, language-specific processing in Llama-3.1, as well as the other tested models, demonstrates a strong concentration in layers 17-28, peaking at layers 27-28. Earlier Layers 0-15 show relatively minimal language-specific activity, with only modest peaks around layers 3-4 and 8-9. This contrasts with previous findings by Tang et al. (2024), Kojima et al. (2024), and Zhao et al. (2024), who report both early- and late-layer language neuron specialization.
Individual language distribution analysis (see Appendices E and F) reveals typologically coherent clustering, with Germanic languages (German (de), Dutch (nl), Afrikaans (af), Danish (da), Norwegian (no), Swedish (sv)) showing similar language neuron distribution patterns concentrated around layers 24-26. Tibetan (bo) and Chinese (zh) (the only Sino-Tibetan languages we tested) demonstrate the most concentrated deep-layer activation, while Romance languages display more varied patterns. These patterns are observed across all tested models.
Figure 3: Layer-wise distribution of language-specific neurons in Llama-3.1 across 21 languages. Similar patterns for individual languages and other models are shown in Appendix E.
Language Neuron Overlap: The language neuron overlap matrices in Figure 2 and Appendices E and F show that language families exhibit internal cohesion with substantial neuron overlap, as also observed by Tan et al. (2024). In LLama-3.1, for instance, Germanic languages show particularly high overlap values (e.g., Swedish (sv) and Danish (da): 290 neurons representing 76.3% and 74.7% of their respective language-specific neurons; Dutch (nl) and Afrikaans (af): 250 neurons representing 76.5% and 59.0% respectively), possibly confirming that the similar activation patterns observed in Figure 13 reflect shared neurons rather than coincidental distributions.
Non-Latin script languages display notably higher neuron counts, with Chinese (zh*: 681), Tibetan (bo*: 1492), and Hindi (hi*: 781) having substantially more language-specific neurons than their Latin-script counterparts. Interestingly, these languages show minimal overlap with Latin-script languages, suggesting that orthographic complexity requires dedicated, non-transferable neural specialization. This is consistent across all models.
Per-layer Language Analysis: For more English-centric models (see Appendix G) such as Llama-3.1 and Mistral-Nemo, we observe that the target language tokens predominantly emerge in the final layers (25–30), aligning with the concentration of language-specific neurons in those layers. English tokens slightly begin to appear in the initial layers and persist through to the output. This pattern partly aligns with findings by Wendler et al. (2024) and Schut et al. (2025). Language entropy in these models increases across layers: the model appears confident in its language prediction in early layers, but entropy rises from around layer 20 onward, suggesting that the decision about which language to generate is made in the final layers.
Figure 4: Predicted probabilities of target languages for Llama-3.1 using logit lens outputs from each layer and FastText for language identification. Results for other models are shown in Appendix G.
In contrast, more multilingual models such as Aya-Expanse-8B and 32B exhibit English token presence across nearly all layers, while the target language is also weakly activated throughout and peaks near the end. These models show an opposite entropy trend: entropy starts high, indicating early uncertainty about the generation language, and gradually decreases toward the final layers, where the model settles on a specific language.
4 Language Neuron Manipulation
4.1 Method
We adopt a manipulation approach inspired by Tang et al. (2024), but instead of replacing neuron activations with fixed values, as done in Tang et al. (2024) (using mean activations) and Kojima et al. (2024) (using median activations), we perform an additive intervention. This additive strategy is less destructive and allows the model to retain more of its original contextual dynamics while still shifting its behavior toward the target language.
Let $\mathbf{a}\in\mathbb{R}^{L\times D}$ be the post-activation tensor after the gated projection layer for a single sequence, where $L$ is the sequence length, and $D$ is the hidden size. Let $\mathcal{I}\subset\{1,\dots,D\}$ be the set of neuron indices identified as language-specific. For each $i\in\mathcal{I}$, we define a boost value $b_{i}$ corresponding to the average activation of neuron $i$ when processing that language.
We apply the following update to the activation tensor:
$\mathbf{a}_{:,i}\leftarrow\mathbf{a}_{:,i}+b_{i}\quad\text{for all }i\in\mathcal{I}$
This additive manipulation shifts the activation patterns toward those typical of the desired language, while preserving the original contextual information encoded in the activations.
Model Intervention Strategy Top 1% Top 2% Top 3% Top 4% Top 5%
Llama-3.1 Additive Activate 30.54% 36.62% 40.66% 42.67% 43.31%
Additive Deactivate + Activate 42.29% 47.99% 52.34% 54.65% 55.4%
Replacement Activate 23.09% 27.85% 31.55% 32.35% 32.8%
Replacement Deactivate + Activate 35.71% 38.85% 42.78% 42.97% 42.78%
DiffMean Activate 28.45% 32.16% 35.52% 37.18% 37.49%
DiffMean Deactivate + Activate 40.51% 44.48% 47.96% 50.15% 51.17%
Table 1: Overall success rates (%) of language forcing for Llama-3.1-8B using three intervention types and two manipulation strategies across different top-k% neuron thresholds.
4.2 Tasks and Data
We evaluate our neuron selection and manipulation method across five multilingual tasks to test its effectiveness in both generative and classification settings. For each task, we experiment with five different fractions of language-specific neurons.
First, we design a novel controlled language-forcing task using six simple questions ("How are you today?", "What is your name?", "What year is it now?", "What is your favorite color?", "What is the weather like?", "Where are you from?") translated into 21 languages using Google Translate Wu et al. (2016). For each input question, we deactivate the neurons associated with the source language by setting their activations to zero or a small negative value (see Appendix H) and activate the neurons associated with a desired target language. We then use FastText (Joulin et al., 2016) to identify the language of the generated output. This setup evaluates the model’s ability to override the input language and generate responses in the specified target language.
We further evaluate our neuron manipulation method across four multilingual downstream tasks covering both generation and classification outputs. For generation, we use FLORES-200 (Costa-Jussà et al., 2022) for machine translation, activating only target-language neurons during decoding to test whether neuron-level steering can enforce correct-language outputs, and XQuAD (Artetxe et al., 2020) for extractive question answering. For classification, we use XNLI (Conneau et al., 2018b) for natural language inference and Belebele (Bandarkar et al., 2024) for multiple-choice reading comprehension. We vary the ratio of manipulated neurons to assess how language-specific activation impacts multilingual performance. All evaluations are performed in a zero-shot setting, with prompts and generation hyperparameters detailed in Appendix A.
Additionally, we investigate whether models exhibit a "fallback" mechanism in language selection during generation. To this end, we use a set of 70 English questions from Vicuna Chiang et al. (2023) and progressively deactivate neurons associated with high-resource languages one by one.
4.3 Results
Language Forcing: For the language forcing task, we compare two neuron manipulation strategies: (1) deactivating neurons associated with the source language and simultaneously activating neurons for the target language, and (2) only activating the neurons corresponding to the desired target language. Our results show that the first approach more reliably steers the model to generate output in the target language (Table 1 and Appendix H). We additionally compare our additive language arithmetic intervention with the simple replacement methods Tang et al. (2024); Kojima et al. (2024) and the widely used DiffMean approach Marks and Tegmark (2023); Panickssery et al. (2023) [^2] For DiffMean, we construct the steering vector as the difference between the mean activation for the target language (positive class) and that of all 20 other languages (negative class). (see Table 1 and Appendix H). We find that the additive method performs best, likely because it is less disruptive to the model’s internal representations.
Figure 1 illustrates the effects of this manipulation across 21 typologically diverse languages in Llama-3.1-8B with the top 5% neuron ratio. The results indicate that higher-resourced languages such as Chinese (zh), Korean (ko), Japanese (ja), German (de), etc. are more susceptible to successful language forcing. In contrast, languages with likely lower representation in the pretraining corpus are more difficult to control, showing inconsistent output.
Below, we provide qualitative examples demonstrating successful language forcing, where the model generates output in a specified target language despite the input being in a different source language:
Forcing Japanese Input: Q: Wie geht es dir heute? A: Output: 今日は元気です。
Forcing German Input: Q: Откуда ты? A: Output: Ich komme aus Deutschland.
Forcing Korean Input: Q: Comment tu t’appelles? A: Output: 저는 김민수입니다..
Figure 5: FLORES performance changes over the baseline (measured by BLEU score) when activating language-specific neurons for Mistral-Nemo (5%).
Figure 6: QXuAD performance changes over the baseline (measured by F1-score) when activating language-specific neurons for Mistral-Nemo (5%).
Figure 7: BELEBELE changes over the baseline (measured by accuracy score) when activating language-specific neurons for Mistral-Nemo (5%).
Figure 8: XNLI changes over the baseline (measured by accuracy score) when activating language-specific neurons for LLama-3.1 (4%).
Downstream Tasks: We evaluate the impact of activating language-specific neurons during inference on four multilingual benchmarks: FLORES-200 and Belebele using all languages covered in our study, XNLI and XQuAD using all available languages. We conduct systematic evaluations on Llama-3.1 and Mistral-Nemo across five neuron ratios. Selected heatmap results are shown in Figures 8, 8, 8, and 8.
Overall, we observe modest improvements (up to 10–20%) or neutral effects for most language pairs when activating neurons from various languages, particularly within the same language family, in contrast to the negative transfer effects reported by Mondal et al. (2025). For instance, activating Romance or Germanic language neurons consistently benefits other languages from the same families across all tasks. Task type influences cross-linguistic effects: translation and question answering show more consistent positive transfer, while machine comprehension and natural language inference exhibit mixed results, including occasional negative transfer from distant languages. Activating the actual target language typically improves performance across all tasks and languages. Interestingly, some Romance and Germanic languages benefit even from activating Sino-Tibetan and Other language neurons–likely due to the larger number of neurons in those groups and their partial overlap with the other families at current activation ratios. Finally, Hindi and Urdu show disproportionate gains from nearly all language activations, possibly due to polysemantic neurons–units encoding multiple concepts (Elhage et al., 2022).
Language "Fallbacks": By progressively deactivating high-resource language neurons in Llama-3.1-8B (see Appendix I), we observe that the model exhibits internal "fallback" strategies for language generation. When English (en) neurons are deactivated, the model defaults to Spanish (es), French (fr), and Russian (ru) in a few cases. Deactivating English and French shifts responses primarily to Russian (ru), with some in Italian (it), Spanish (es), and Arabic (ar). Further removal of Latin-script languages–Spanish, Italian, and Portuguese (pt)–promotes generation in Russian, Arabic, and Chinese. Deactivating German, Russian, and Arabic leads to "fallback" responses in Thai, Vietnamese, and Chinese. Notably, some English, Spanish, and French responses persist despite deactivation, reflecting the model’s strong priors for these languages. Similar patterns are observed for Mistral-Nemo, but with a slightly different language hierarchy. We leave a more detailed investigation of these dynamics to future work.
5 Discussion
5.1 Identification Insights
The layer-wise distribution of language-specific neurons in Llama-3.1, Mistral Nemo, and Aya-Expanse partially aligns with prior findings Tang et al. (2024); Kojima et al. (2024); Zhao et al. (2024). We observe that these neurons are predominantly concentrated in the later layers, whereas Tang et al. (2024) and Kojima et al. (2024) report a dual-peaked distribution in older models, with concentrations in both early and late layers. Notably, Tang et al. (2024) also show that the more multilingual BLOOM model Le Scao et al. (2023) exhibits a similar late-layer concentration, a trend further supported by Mondal et al. (2025) with recent multilingual models. This divergence likely reflects architectural differences, or variations in training data composition. Interestingly, Chou et al. (2025) demonstrate that modifying a small number of SAE features from mid-to-late layers enables high-accuracy language steering for high-resource languages, while Andrylie et al. (2025) similarly find that their SAE-LAPE method identifies language-specific features concentrated in the same layers. These converging results, obtained through entirely different methodologies, reinforce our observation that language-specific representations are predominantly localized in the mid-to-late transformer layers of modern multilingual LLMs.
Layer-wise logit lens analysis complements these findings: in all tested models, English token probabilities emerge early on and persist, while target language tokens appear more sharply in the very final layers, which is in line with Wendler et al. (2024) and Schut et al. (2025). Entropy trends in more English-centric models rise toward the end, suggesting that language generation decisions are finalized in the later layers. In terms of cross-linguistic neuron overlap, we observe clear clustering patterns based on language families. This supports findings from Tan et al. (2024), who report that related languages share substantial neural components while maintaining distinct processing features.
5.2 Manipulation Insights
Our findings show that neuron-level interventions can reliably steer multilingual model behavior, with our additive language arithmetic method outperforming activation replacement Tang et al. (2024); Kojima et al. (2024) and DiffMean-based approaches Marks and Tegmark (2023). Activating language-specific neurons improves performance across both generative and discriminative tasks, with the strongest gains observed when the activated language matches the target or is typologically related, yielding positive transfer within language families, but more limited benefits for structurally distant languages. Interestingly, Hindi and Urdu show disproportionate improvements across many interventions, potentially due to neurons encoding multiple overlapping functions (Elhage et al., 2022). Moreover, compared to Mondal et al. (2025), who report limited improvements from activation replacement or language neuron fine-tuning on tasks like XNLI Conneau et al. (2018a) and XQuAD Artetxe et al. (2019), our method achieves consistent gains due to three key differences. First, we evaluate across a broader and more typologically diverse set of languages. Second, our additive interventions using mean activation vectors preserve more of the original representational context. Third, all evaluations are conducted in a strict zero-shot setting, without fine-tuning on downstream tasks in any source language.
The effects are especially pronounced in our language forcing experiments. Deactivating neurons for the input language while simultaneously activating neurons for a desired target language significantly increases the likelihood of generating output in the target language–even when the input and output languages are typologically distant. These results suggest that target-language neurons contain sufficient signal to override source-language priors, especially for languages better represented in pre-training. In contrast, low-resource languages are harder to control, likely due to weaker or less distinct neural representations resulting from limited pretraining exposure. Importantly, manipulation also reveals internal "fallback" mechanisms: when dominant language neurons like English or French are deactivated, the model switches to secondary high-resource languages such as Russian or Arabic. This "fallback" hierarchy varies slightly across models but consistently exposes the model’s internal language preferences. The persistence of certain languages despite deactivation–particularly English–highlights the robustness of high-priority language priors in LLMs.
6 Conclusion
Our large-scale study demonstrates that language-specific neurons encode meaningful linguistic features that can be identified, analyzed, and manipulated to steer model behavior across languages. By employing a novel additive intervention method, we show that activating these neurons allows for generation language control and improves performance on various multilingual tasks. These findings underscore the potential of neuron-level control as an interpretable mechanism for guiding multilingual model outputs.
Limitations
Our study has several limitations. First, while we demonstrate the effectiveness of neuron-level interventions across multiple tasks, a broader evaluation spanning more diverse downstream settings–such as dialogue, summarization, or code generation–is needed to assess the generality and robustness of our approach. Second, although our interventions often produce outputs in the desired language, the quality and fluency of these forced outputs across different languages–especially lower-resourced ones–remain underexplored and require more rigorous evaluation. Third, our current analysis primarily focuses on isolated neurons with high language specificity. This overlooks the possibility that language representations may be distributed across larger sub-networks or circuits; future work should investigate the interactions and dependencies between such neurons to better understand the network-level mechanisms supporting multilingual processing.
Additionally, due to computational constraints, not all experiments were run across every model in our study. In particular, some analyses (large-scale downstream evaluation) were limited to smaller models–LLama-3.1-8B and Mistral-Nemo-14B. Expanding these evaluations to include larger models could reveal further insights but requires significantly more resources.
Acknowledgments
This research was supported by DisAI - Improving scientific excellence and creativity in combating disinformation with artificial intelligence and language technologies, a Horizon Europe-funded project under GA No. 101079164, and by the German Federal Ministry of Research, Technology and Space (BMFTR) as part of the project TRAILS (01IW24005).
Appendix
Appendix A Prompt Templates
A.1 FLORES-200 Machine Translation
For the machine translation task, we use the following prompt format to instruct the model to translate from a source language into a target language:
Translate this {source_name} sentence into {target_name}: {source_text}. Translation:
Here, {source_name} and {target_name} are replaced by the names of the source and target languages (e.g., French, Hindi), and {source_text} is the input sentence.
A.2 XQuAD Question Answering
For extractive question answering, we prompt the model to find an answer span in a given context:
Answer the question based on the {language_name} context provided. Extract the exact answer from the context.
Context: {context}
Question: {question}
Answer:
Here, {language_name} is the name of the language (e.g., Hindi), and {context}, {question} are the inputs.
A.3 Belebele Machine Comprehension
For the multiple-choice machine comprehension task, we use:
Read the following {language_name} passage and answer the question.
Passage: {passage}
Question: {question}
A. {option_A}
B. {option_B}
C. {option_C}
D. {option_D}
Answer:
Each {option_X} is a possible answer choice, and the model is expected to return one of A, B, C, or D.
A.4 XNLI Natural Language Inference
To classify the relationship between a premise and hypothesis, we use:
Given the following {language_name} premise and hypothesis, determine the relationship between them.
Premise: {premise}
Hypothesis: {hypothesis}
Options:
1. Entailment
2. Neutral
3. Contradiction
Answer:
The model is expected to return one of the listed option numbers or labels.
A.5 Generation Settings
We use the same decoding configuration across all tasks unless specified otherwise. The generation is performed using the following sampling parameters:
SamplingParams(    temperature=0,    repetition_penalty=1.1,    stop_token_ids=[eos_token_id] if eos_token_id is not None else [],    skip_special_tokens=True )
We vary the max_tokens parameter depending on the task:
  • FLORES-200 (Machine Translation): max_tokens = 128
  • XQUAD (Question Answering): max_tokens = 64
  • XNLI (Natural Language Inference): max_tokens = 32
  • BELEBELE (Machine Comprehension): max_tokens = 32
  • Language Forcing Experiments: max_tokens = 256
These settings ensure deterministic generation (due to temperature=0) while reducing repetition and enabling flexible truncation across task types.
Appendix B Selected Languages
Language Code Language Name
bo Tibetan
mt Maltese
it Italian
es Spanish
de German
ja Japanese
ar Arabic
zh Chinese
af Afrikaans
nl Dutch
fr French
pt Portuguese
ru Russian
ko Korean
hi Hindi
tr Turkish
pl Polish
sv Swedish
da Danish
no Norwegian
en English
Table 2: Language codes and corresponding full names
Appendix C Neuron Counts
Lang. 1% 2% 3% 4% 5%
bo 1492 2687 3723 4708 5687
mt 617 1578 2687 3853 5043
it 337 761 1198 1653 2104
es 298 667 1007 1358 1717
de 284 651 1065 1509 1965
ja 638 1084 1503 1887 2286
ar 400 708 1093 1472 1913
zh 681 1150 1564 1929 2286
af 455 1174 1966 2852 3803
nl 363 901 1432 2040 2680
fr 269 615 988 1376 1786
pt 310 698 1072 1449 1844
ru 394 717 985 1248 1546
ko 556 830 1054 1264 1484
hi 781 1456 2071 2644 3199
tr 476 1031 1681 2307 2916
pl 378 900 1455 2065 2744
sv 420 990 1614 2358 3089
da 431 1055 1771 2557 3344
no 433 1041 1743 2499 3268
en 61 88 116 140 151
Table 3: Number of language-specific neurons at different top-k thresholds in Llama-3.1.
Lang. 1% 2% 3% 4% 5%
bo 1153 1941 2547 3059 3554
mt 250 717 1283 1913 2552
it 155 343 539 766 969
es 154 343 528 730 907
de 130 291 460 656 859
ja 384 770 1113 1446 1808
ar 275 499 703 912 1131
zh 346 717 1076 1380 1723
af 219 567 1012 1549 2115
nl 184 458 791 1207 1646
fr 156 323 475 650 812
pt 175 388 589 799 1001
ru 213 387 520 653 764
ko 335 623 852 1079 1299
hi 310 619 924 1193 1443
tr 213 586 1027 1510 1982
pl 186 436 752 1106 1472
sv 160 451 807 1211 1637
da 171 486 886 1335 1812
no 148 464 836 1237 1643
en 32 55 78 100 117
Table 4: Number of language-specific neurons at different top-k thresholds in Mistral-Nemo-Base-2407.
Lang. 1% 2% 3% 4% 5%
bo 1636 3343 4947 6551 8085
mt 433 1169 2108 3079 4150
it 231 591 981 1393 1850
es 277 621 1016 1388 1745
de 211 509 874 1245 1674
ja 636 1171 1580 1983 2393
ar 383 662 962 1251 1566
zh 702 1348 1933 2512 3030
af 296 781 1440 2181 2925
nl 181 485 859 1243 1633
fr 211 489 842 1229 1636
pt 280 621 1007 1408 1811
ru 381 708 1052 1356 1671
ko 597 1051 1405 1763 2110
hi 459 880 1269 1668 2063
tr 322 734 1126 1512 1927
pl 260 633 1003 1364 1746
sv 282 837 1595 2437 3331
da 303 886 1688 2628 3569
no 290 866 1672 2576 3496
en 55 80 115 141 169
Table 5: Number of language-specific neurons at different top-k thresholds in Aya-Expanse-8B.
Lang. 1% 2% 3% 4% 5%
bo 1107 2241 3314 4403 5541
mt 568 1431 2403 3427 4461
it 353 758 1192 1629 2076
es 344 711 1100 1483 1805
de 240 631 1105 1557 1986
ja 676 1249 1703 2152 2563
ar 361 643 902 1163 1436
zh 656 1306 1920 2507 3032
af 364 978 1672 2368 3021
nl 235 607 990 1371 1691
fr 280 593 943 1332 1717
pt 362 709 1134 1548 1940
ru 355 757 1138 1514 1860
ko 549 1000 1383 1746 2091
hi 498 956 1296 1604 1899
tr 331 728 1138 1524 1902
pl 276 656 1036 1467 1823
sv 322 897 1585 2273 2925
da 343 980 1739 2488 3162
no 329 968 1705 2448 3102
en 51 85 122 148 171
Table 6: Number of language-specific neurons at different top-k thresholds in Aya-Expanse-32B.
Appendix D Downstream Baseline Results
Model af ar bo da de es fr hi it ja ko mt nl no pl pt ru sv tr zh
Llama-3.1 0.227 0.049 0.172 0.233 0.176 0.154 0.276 0.113 0.155 0.024 0.040 0.086 0.136 0.163 0.100 0.311 0.166 0.217 0.096 0.102
Mistral-Nemo 0.114 0.022 0.176 0.109 0.140 0.144 0.215 0.023 0.123 0.140 0.022 0.030 0.088 0.076 0.039 0.190 0.104 0.069 0.041 0.136
(a) Machine translation (BLEU scores).
Model af ar bo da de en es fr hi it ja ko mt nl no pl pt ru sv tr zh
Llama-3.1 0.690 0.760 0.316 0.742 0.764 0.847 0.751 0.764 0.670 0.766 0.729 0.767 0.542 0.750 0.744 0.714 0.777 0.799 0.787 0.727 0.779
Mistral-Nemo 0.689 0.751 0.284 0.706 0.792 0.841 0.823 0.790 0.357 0.806 0.749 0.798 0.408 0.733 0.697 0.659 0.783 0.840 0.714 0.734 0.808
(b) Machine comprehension (accuracy).
Model ar bg de el en es fr hi ru sw th tr ur vi zh
Llama-3.1 0.382 0.361 0.365 0.356 0.429 0.358 0.377 0.289 0.375 0.350 0.386 0.397 0.130 0.358 0.379
Mistral-Nemo 0.332 0.333 0.335 0.335 0.338 0.333 0.349 0.331 0.335 0.332 0.334 0.335 0.327 0.334 0.334
(c) Natural language inference (accuracy).
Model ar de el en es hi ro ru th tr vi zh
Llama-3.1 0.410 0.462 0.422 0.616 0.515 0.460 0.542 0.416 0.459 0.511 0.606 0.355
Mistral-Nemo 0.281 0.443 0.258 0.585 0.556 0.387 0.451 0.362 0.379 0.452 0.444 0.276
(d) Question answering (F1 score).
Table 7: Baseline multilingual evaluation results for translation, comprehension, NLI, and QA.
Appendix E Language Neuron Distributions
Figure 9: Distribution of identified language-specific neurons across Mistral-Nemo layers for all 21 evaluated languages. The neuron distributions for individual languages are further in the Appendix.
Figure 10: Distribution of individual language-specific neurons across Llama-3.1 layers for all 21 languages.
Figure 11: Distribution of individual language-specific neurons across Mistral-Nemo layers for all 21 languages.
Figure 12: Overlap of language-specific neurons between individual languages and language families in Mistral-Nemo when considering top 1% of neurons as potentially language-specific. Diagonal values indicate the number of language-specific neurons for each language; off-diagonal values indicate the number of overlapping neurons. Asterisks denote languages with non-Latin scripts.
Figure 13: Distribution of individual language-specific neurons across Llama-3.1-8B layers for all 21 languages.
Figure 14: Distribution of individual language-specific neurons across Mistral-Nemo layers for all 21 languages.
Appendix F Additional Results for Aya Models
Figure 15: Distribution of identified language-specific neurons across Aya-Expanse-8B layers for all 21 evaluated languages.
Figure 16: Distribution of individual language-specific neurons across Aya-Expanse-8B layers for all 21 languages.
Figure 17: Distribution of identified language-specific neurons across Aya-Expanse-32B layers for all 21 evaluated languages. The neuron distributions for individual languages are further in the Appendix.
Figure 18: Distribution of individual language-specific neurons across Aya-Expanse-32B layers for all 21 languages.
Figure 19: Distribution of individual language-specific neurons across Aya-Expanse–8B layers for all 21 languages.
Figure 20: Distribution of individual language-specific neurons across Aya-Expanse-32B layers for all 21 languages.
Figure 21: Overlap of language-specific neurons between individual languages and language families in Aya-Expanse-8B and 32B when considering top 1% of neurons as potentially language-specific. Diagonal values indicate the number of language-specific neurons for each language; off-diagonal values indicate the number of overlapping neurons. Asterisks denote languages with non-Latin scripts.
Appendix G Output Analysis with Logit Lens
(a) Probabilities of the target languages.
(b) Probabilities of English.
(c) Evolution of target language probability.
(d) Heatmap of English probabilities.
Figure 22: Peeking into language prediction across layers using the logit lens for Llama-3.1-8B.
(a) Probabilities of the target languages.
(b) Probabilities of English.
(c) Evolution of target language probability.
(d) Heatmap of English probabilities.
Figure 23: Peeking into language prediction across layers using the logit lens for Mistral-Nemo.
(a) Probabilities of the target languages.
(b) Probabilities of English.
(c) Evolution of target language probability.
(d) Heatmap of English probabilities.
Figure 24: Peeking into language prediction across layers using the logit lens for Aya-Expanse-8B.
(a) Probabilities of the target languages.
(b) Probabilities of English.
(c) Evolution of target language probability.
(d) Heatmap of English probabilities.
Figure 25: Peeking into language prediction across layers using the logit lens for Aya-Expanse-32B.
Appendix H Language Forcing
(a) Activate-only - Top 1%
(b) Activate-only - Top 2%
(c) Activate-only - Top 3%
(d) Deactivate-Activate - Top 1%
(e) Deactivate-Activate - Top 2%
(f) Deactivate-Activate - Top 3%
Figure 26: Page 1 - Language forcing results across 5 neuron ratios for two manipulation strategies - activate-only and deactivate-activate for Llama-3.1-8B. For deactivation, we set the neurons to 0.
(a) Activate-only - Top 4%
(b) Activate-only - Top 5%
(c) Deactivate-Activate - Top 4%
(d) Deactivate-Activate - Top 5%
Figure 27: Page 2 - Language forcing results across 5 neuron ratios for two manipulation strategies - activate-only and deactivate-activate for Llama-3.1-8B. For deactivation, we set the neurons to 0.
Appendix I Language "Fallbacks"
(a) Top 4% neurons
(b) Top 5% neurons
Figure 28: Progressive deactivation of language-specific neurons for high-resource languages for Llama-3.1-8B. We set the language neurons to -1 for deactivation in this scenario; 0 does not produce the required effects.
(a) Top 4% neurons
(b) Top 5% neurons
Figure 29: Progressive deactivation of language-specific neurons for high-resource languages for Mistral-Nemo. We set the language neurons to -2.5 for deactivation in this scenario; values from 0 to -2 do not produce the required effects.